OpenStudy (anonymous):

Choose the correct rewriting of this algebraic fraction and give the exclusion represented by your choice. 4y (8y + 24) –1 = y ≠

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (campbell_st):

you may want to retype the question, as there doesn't appear to be a fraction

OpenStudy (anonymous):

Answer choices for #1: 2y/2y+3 or y/(2)(y+3) #2: -3 or 3

OpenStudy (anonymous):

Is that better? @campbell_st

OpenStudy (campbell_st):

takes a fraction... so looking at the 2 terms in the denominator... do 2y and 3 have a common factor...?

OpenStudy (anonymous):

No

OpenStudy (campbell_st):

can I check is the fraction 1. \[\frac{4y}{8y + 24} -1 \]

OpenStudy (anonymous):

What?

OpenStudy (campbell_st):

well the original fraction you posted has -1 at the end...

OpenStudy (anonymous):

No, that's an exponent.

OpenStudy (campbell_st):

ok... so its \[4y (8y + 24)^{-1}\]

OpenStudy (campbell_st):

which can be written as \[\frac{4y}{8y + 24}\] 4 is a common factor... so divide every term by 4 what do you get...?

OpenStudy (anonymous):

y/2y+6

OpenStudy (campbell_st):

great.... and is ther a common factor in the denominator..?

OpenStudy (anonymous):

2?

OpenStudy (campbell_st):

great so then the factored form is a little weird but its the 2nd of your choices... so looking at y + 3 what value of y makes the expression zero..?

OpenStudy (anonymous):

Dude I have no idea.

OpenStudy (campbell_st):

well solve this Dudette y + 3 = 0

OpenStudy (campbell_st):

the answer you get will be the restriction for the denominator

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

correct... so now you have the solution.... restrictions in fractions are the value(s) that make the denominator zero.... so looking at the original question you needed to solve 8y + 24 = 0 its the same answer as the simplified version.

OpenStudy (anonymous):

That doesn't make sense

OpenStudy (anonymous):

@superdavesuper

OpenStudy (campbell_st):

ok.... tell be what happens when you enter this into a calculator \[2 \div 0 =\]

OpenStudy (anonymous):

0

OpenStudy (campbell_st):

no actually pick up a calculator and do it

OpenStudy (anonymous):

Lol no. I don't have time for this. You're making me do extra work that is totally unnecessary.

OpenStudy (campbell_st):

you answer is to the problem \[0 \div 2 = 0\]

OpenStudy (campbell_st):

well it is necessary as its fundamental to the questions you are asking.

OpenStudy (anonymous):

Not really. I'm not gonna do extra work when I have enough as it is.

OpenStudy (campbell_st):

is a simple task... use a calculator http://web2.0calc.com/ here is an online on... tell me what happens with the calculation \[2 \div 0 =\]

OpenStudy (anonymous):

@nehemiah2

OpenStudy (anonymous):

Nah

OpenStudy (ribhu):

y not equal to -3

OpenStudy (anonymous):

@ribhu English, please.

OpenStudy (ribhu):

y would not be equal to -3

OpenStudy (anonymous):

Soooo what's that mean?

OpenStudy (campbell_st):

that's a solution you found a lot earlier..

OpenStudy (ribhu):

what do u want exactly?

OpenStudy (anonymous):

Okay, what about the other one then? @ribhu obviously I want the answer.

OpenStudy (campbell_st):

you have found both answers... you just need to go back and read thinks... remember Open study is about helping your understanding... and not for providing answers...

OpenStudy (anonymous):

@nehemiah2

OpenStudy (anonymous):

I'm not gonna get on Facebook, @ribhu